Charlie Company

The Balance Sheet and Income Statement of Charlie Company as of and for the twelve months ended December 31, 2018 shows the following information (in thousands):

Opening (Prior Period) and Ending balances

Accounts Receivable Opening $4,155, Ending. $5,155

Accounts Payable Opening $5,200, Ending $5,750

Inventory Opening $2,150, Ending 4,150

Sales on Credit (12months) $42,450

Cost of Goods Sold (COGS) (12 months) 31,500

Utilizing the information in question 1 above for Charlie Company in 2018 the number of days in receivable were _____ (round to closest whole number of days)?

A 22 days

B 28 days

C 40 days

D 36 days
Answer #1
Average Receivables= ( beginning Receivables+ ending Receivables)/2
Average Receivables = (4155+5155)/2
Average Receivables = 4655
Receivables turnover = Credit sales/receivables
Receivables turnover = 42450/4655
Receivables turnover = 9.12
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/9.12
days of sales outstanding = 40.02